Calculate the molecular number of 1.22 mole of H2SO4
1
Expert's answer
2017-05-08T08:30:09-0400
N = NA × n = 6.62×10^23 × 1.22 = 8.08×10^23 molecules

Answer:
N = 8.08×10^23 molecules